OpenStudy (anonymous):

If the circumference of a circle of radius r inches is equal to the perimeter of a square with aside length of s inches so r/s=?

OpenStudy (anonymous):

Do you have the formula for the circumference of a circle?

OpenStudy (anonymous):

Yes 2pi\[^{2}\]

OpenStudy (anonymous):

almost... it's 2 π r like 2 times pi times the radius.

OpenStudy (anonymous):

And what is the perimeter of the square with sides that are "s" inches each? You won't have a number, but just add up each of the 4 sides with length "s"... you get "4s" Are you with me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

This problem says the the circle's circumference is equal to the square's perimeter. Let's write an equation that shows they are equal...

OpenStudy (anonymous):

2πr = 4s Does that make sense? You have circumference on the left, and perimeter on the right.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so the last thing the problem asks is, what is r / s ? You need to take that equation and shift things around until you have r/s on one side of the equation... then it equals the stuff on the other side. 2πr = 4s ----->>>> r/s = ???

OpenStudy (anonymous):

im confused what to do when i get pi r divided by s =2

OpenStudy (anonymous):

Try to do the next step... you might divide both sides by "s".... that way you would at least have an "r/s" with some other stuff on the left side... it's a decent start.

OpenStudy (anonymous):

is that what you get? π r/s = 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You're almost done :) Now divide both sides by π to leave r/s all alone on the left side

OpenStudy (anonymous):

r/s = 2/π If you need to show it as numbers, hopefully they told you how to estimate "pi"... like if they said "estimate pi as 3.14", then you can do 2/3.14 as the solution. Otherwise, just leave it as 2/π

OpenStudy (anonymous):

So you don't really know or even need to know what "r" and "s" were... you just found their ratio, just like the problem asked... It was just a matter of writing the formulas for each, setting them equal, and then simplifying.